Please help me solve this equation:
the square root of 7a+1=a+1

Answer by edjones(8007)

sqrt(7a+1)=a+1 I assume that this is the problem.
7a+1=a^2+2a+1 Square each side.
a^2-5a=0
a(a-5)=0
a=0, a=5
.
Ed